# Combining Crown Structures for Vulnerability Measures

### v1

@@ Abstract
Over the past decades, various metrics have emerged in graph theory to grasp the complex nature of network vulnerability. In this paper, we study two specific measures: (weighted) vertex integrity (wVI) and (weighted) component order connectivity (wCOC). These measures not only evaluate the number of vertices required to decompose a graph into fragments, but also take into account the size of the largest remaining component. The main focus of our paper is on kernelization algorithms tailored to both measures. We capitalize on the structural attributes inherent in different crown decompositions, strategically combining them to introduce novel kernelization algorithms that advance the current state of the field. In particular, we extend the scope of the balanced crown decomposition provided by Casel et al. [DBLP:conf/esa/Casel0INZ21] and expand the applicability of crown decomposition techniques.
In summary, we improve the vertex kernel of VI from $p^{3}$ to $p^{2}$, and of wVI from $p^{3}$ to $3(p^{2}+p^{1.5}p_{\ell})$, where $p_{\ell}<p$ represents the weight of the heaviest component after removing a solution. For wCOC we improve the vertex kernel from $\mathcal{O}(k^{2}W+kW^{2})$ to $3\mu(k+\sqrt{\mu}W)$, where $\mu=\max(k,W)$. We also give a combinatorial algorithm that provides a $2kW$ vertex kernel in FPT-runtime when parameterized by $r$, where $r\leq k$ is the size of a maximum $(W+1)$-packing. We further show that the algorithm computing the $2kW$ vertex kernel for COC can be transformed into a polynomial algorithm for two special cases, namely when $W=1$, which corresponds to the well-known vertex cover problem, and for claw-free graphs. In particular, we show a new way to obtain a $2k$ vertex kernel (or to obtain a 2-approximation) for the vertex cover problem by only using crown structures.
@@ 1 Introduction
In the study of graph theory different scales have emerged over the past decades to capture the complex nature of network vulnerability. While the main focus is on the connectivity of vertices and edges, subtle aspects of vulnerability such as the number of resulting components, the size distribution of the remaining components, and the disparity between them are becoming increasingly interesting. Our focus in this study is on two specific ways to measure vulnerability: (weighted) vertex integrity and (weighted) component order connectivity. These measures not only evaluate the number of vertices required to decompose a graph into fragments, but also take into account the size of the largest remaining component. Incorporating these aspects provides a more comprehensive understanding of network resilience.
Informally, vertex integrity (VI) models finding the right balance between removing vertices and keeping small connected parts of a graph. More formally, given a graph $G=(V,E)$ and a number $p\in\mathds{N}$, VI tries to find a set of vertices $S\subseteq V$ such that $|S|$ plus the size of the biggest component in the remaining graph is at most $p$. In the vertex weighted version (wVI), the goal is bounding the total weight of the removed vertices plus the weight of the heaviest component by $p$. This problem was introduced by Barefoot et al. [barefoot1987vulnerability] as a way to measure vulnerability of communication networks. Recently, it has drawn the interest of the parameterized complexity community due to its status as a natural parameter that renders numerous NP-hard problems amenable to fixed parameter tractability (FPT). This means that for these problems, solutions can be computed within a time frame represented by $f(p)\cdot n^{\mathcal{O}(1)}$, where $f$ is a computable function [DBLP:conf/isaac/LampisM21]. It is interesting to see how vertex integrity relates to other well-known measures of network structure. It imposes greater constraints compared to metrics such as treedepth, treewidth, or pathwidth, as the vertex integrity of a graph serves as an upper bound for these parameters. However, it encompasses a wider range of scenarios compared to vertex cover, where a vertex cover of a graph is an upper bound for its vertex integrity. This makes it a key in understanding how to efficiently solve problems in the world of network analysis.
@@
The focus of the paper is on kernelization algorithms tailored for both weighted and unweighted versions of VI and COC when parameterized by $p$ and $k+W$, respectively. Kernelization algorithms can be thought of as formalized preprocessing techniques aimed to reduce optimization problems. Of particular interest in this work are crown decompositions, which are generally used as established structures for safe instance reduction - where "safe" means that there is no negative impact on the corresponding optimization problem. Essentially, a crown decomposition partitions the vertex set into distinct components: crown, head, and body. Here, the head acts as a separator between the crown and the body. This structural arrangement becomes useful when specific relationships between the head and the crown are required. Such relationships ultimately enable us to shrink instances by eliminating these designated parts from the graph. The properties of this structural layout, coupled with its existence guarantees depending on the instance size, enable the development of efficient kernelization algorithms for different problem domains. Notably, crown decompositions have also recently found utility in approximation algorithms for graph packing and partitioning problems [DBLP:conf/esa/Casel0INZ21]. For further exploration of crown decompositions, including their variations and applications, we recommend the comprehensive survey paper by Jacob et al. [DBLP:journals/algorithms/JacobMR23].
@@
Recently, the vertex integrity problem (VI) was extensively studied as a structural graph parameter [DBLP:conf/esa/BentertHK23, DBLP:journals/tcs/GimaHKKO22, DBLP:journals/algorithmica/GimaO24, DBLP:conf/isaac/LampisM21]. Gima et al. [DBLP:conf/walcom/GimaHKM0O24] conducted a systematic investigation into structural parameterizations computing the VI and wVI which was further extended by Hanaka et al. [DBLP:journals/corr/abs-2402-09971]. Additionally, there are notable results concerning special graph classes [DBLP:journals/dam/BaggaBGLP92, clark1987computational, DBLP:journals/algorithmica/DrangeDH16, DBLP:journals/dam/KratschKM97, DBLP:journals/ijcm/LiZZ08]. In our context, regarding related work on VI and wVI, Fellows and Stueckle presented an algorithm that solves the problem in $\mathcal{O}(p^{3p}n)$ [fellows1989immersion]. This was subsequently improved by Drange et al. [DBLP:journals/algorithmica/DrangeDH16], even for the weighted case, to $\mathcal{O}(p^{p+1}n)$. In the same paper, they presented the first vertex-kernel of size $p^{3}$ for both VI and wVI.
@@
We turn now our attention to the related work of COC and wCOC. It is unlikely that kernelization algorithms can be achieved by considering $k$ or $W$ alone in polynomial time. Indeed, $W=1$ corresponds to a hard problem, which shows that $W$ (alone) is not a suitable parameter. For the parameter $k$, the problem is $W[1]$-hard even when restricted to split graphs [DBLP:journals/algorithmica/DrangeDH16]. These lower bounds lead to the study of parameterization by $k+W$. The best known algorithm with respect to these parameters finds a solution in time $n^{O(1)}\cdot 2^{\mathcal{O}(\log(W)\cdot k)}$ [DBLP:journals/algorithmica/DrangeDH16]. Unless the exponential time hypothesis fails, the authors prove that this running time is tight up to constant factors, i.e., there is no algorithm that solves the problem in time $n^{\mathcal{O}(1)}\cdot 2^{o(\log(W)\cdot\text{OPT})}$. The best known approximation algorithm has a multiplicative gap guarantee of $\mathcal{O}(\log(W))$ to the optimal solution with a running time of $n^{\mathcal{O}(1)}\cdot 2^{\mathcal{O}(W)}$ [DBLP:journals/corr/Lee16c]. They showed that the superpolynomial dependence on $W$ may be needed to achieve a polylogarithmic approximation. Using this algorithm as a subroutine, the vertex integrity can be approximated within a factor of $\mathcal{O}(\log(\text{OPT}))$, where OPT is the vertex integrity.
@@
Regarding kernelization algorithms, there is a sequence of results which successively improve the vertex-kernel of COC. The first results came from Chen et al. [DBLP:journals/tcs/ChenFSWY19] and Drange et al. [DBLP:journals/algorithmica/DrangeDH16], who provided a $\mathcal{O}(kW^{3})$ and  $\mathcal{O}(k^{2}W+kW^{2})$ kernel, respectively. The result of Drange et al. also holds for wCOC and is the only result for this case. This was improved simultaneously by Xiao [DBLP:journals/jcss/Xiao17a] as well as by Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16] to a $\mathcal{O}(kW^{2})$ kernel. These works also provide the first $\mathcal{O}(kW)$ kernels, but with different constants and running times. Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16] present a $2kW$ kernel in a running time of $n^{\mathcal{O}(W)}$ by using linear programming (LP) methods with an exponential number of constraints (cf. LABEL:fig::LPWSep). The runtime can be improved to $2^{\mathcal{O}(W)}\cdot n^{\mathcal{O}(1)}$ as already mentioned in the book of Fomin et al. [fomin2019kernelization] (Section 6.4.2). Roughly speaking, the idea is to use the ellipsoid method with separation oracles to solve the linear program, where the separation oracle uses a method called color coding to find violated constraints that makes it tractable in $W$. Note that if $W$ is a constant this $2kW$ kernel is a polynomial time kernel improving on some previous results. This includes for instance the improvement of the $5k$ kernel provided by Xiao and Kou [DBLP:conf/tamc/XiaoK17] to a $4k$ kernel for the well-studied $P_{2}$-covering problem, where a $P_{2}$ is a path with 2 edges. The first linear kernel in both parameters in polynomial time, i.e. an $\mathcal{O}(kW)$ vertex kernel, is presented by Xiao [DBLP:journals/jcss/Xiao17a], who provide a $9kW$ vertex kernel. Finally, this was improved recently by Casel et al. [DBLP:conf/esa/Casel0INZ21] to a $3kW$ vertex kernel, which also holds for a more general setting. Namely, to find $k$ vertices in a vertex weighted graph such that after their removal each component weighs at most $W$. Note that the weights of the separator, i.e. the chosen $k$ vertices, play no role in this problem compared to wCOC. With the exception of the $\mathcal{O}(k^{2}W+kW^{2})$ vertex kernel of Drange et al. [DBLP:journals/algorithmica/DrangeDH16], all achieved vertex kernels essentially use crown structures.
Our Contribution:
@@
The provided running times are based on an input graph $G=(V,E)$. We improve the vertex kernel for VI from $p^{3}$ to $3p^{2}$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$. For wVI, we improve it to $3(p^{2}+p^{1.5}p_{\ell})$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$, where $p_{\ell}<p$ represents the weight of the heaviest component after removing a solution.
@@
To better explain the results of COC and wCOC, we introduce the problem of a maximum $\lambda$-packing for $\lambda\in\mathds{N}$. Given a (vertex weighted) graph $G$, this problem aims to maximize the number of disjoint connected subgraphs, each of size (or weight) at least $\lambda$. It is worth noting that the size of a maximum $(W+1)$-packing serves as a lower bound on $k$, since each element in the packing contains at least one vertex of a solution of COC (or wCOC) — in terms of linear programming, it is the dual of COC (cf. LABEL:fig::LPWSep).
@@
For wCOC we improve the vertex kernel of $\mathcal{O}(k^{2}W+kW^{2})$ to $3\mu(k+\sqrt{\mu}W)$ in time $\mathcal{O}\left(r^{2}k|V|^{2}|E|\right)$, where $\mu=\max(k,W)$ and $r\leq|V|$ is the size of a maximum $(W+1)$-packing. For the unweighted version, we provide a $2kW$ vertex kernel in an FPT-runtime of $\mathcal{O}(r^{3}|V||E|\cdot r^{\min(3r,k)})$, where $r\leq k$ is the size of a maximum $(W+1)$-packing. In particular, for the vertex cover problem (where $W=1$), the kernelization result yields a $2k$ vertex kernel, thereby implying a $2$-approximation. That is, if we consider kernelization for COC within polynomial time to achieve a vertex kernel of size $ckW$ for some constant $c$, the optimal constant achievable is $c=2$, unless the unique games conjecture fails [KHOT2008335]. Comparing this result with the state of the art, disregarding the FPT-runtime aspect, we improve upon the best-known polynomial algorithm, achieving a kernel of size $3kW$ [DBLP:conf/esa/Casel0INZ21]. A $2kW$ vertex kernel is also presented in [DBLP:conf/iwpec/KumarL16], albeit with an exponential runtime using linear programming methods, which, as mentioned, can be enhanced to an FPT-runtime regarding parameter $W$. In contrast, our result is entirely combinatorial and has an FPT-runtime in the parameter of a maximum $(W+1)$-packing $r\leq k$. We further show that the algorithm computing the $2kW$ vertex kernel for COC can be transformed into a polynomial algorithm for two special cases. The first case arises when $W=1$, i.e., for the vertex cover problem. Here, we provide a new method for obtaining a vertex kernel of $2k$ (or obtaining a $2$-approximation) using only crown decompositions. The second special case is for the restriction of COC to claw-free graphs.
@@
Regarding these special cases, until 2017, a $2k$ vertex kernel for $W=1$ was known through crown decompositions, albeit computed using both crown decompositions and linear programming. Previously, only a $3k$ vertex kernel was known using crown decompositions alone. In 2018, Li and Zhu [DBLP:journals/tcs/LiZ18] provided a $2k$ vertex kernel solely based on crown structures. They refined the classical crown decomposition, which possessed an additional property allowing the remaining vertices of the graph to be decomposed into a matching and odd cycles after exhaustively applying the corresponding reduction rule. In contrast, our algorithm identifies the reducible structures, which must exist if the size of the input graph exceeds $2k$. Unfortunately, we were unable to transform the FPT-runtime algorithm into a polynomial-time algorithm in general. Nevertheless, we believe that our insights into the structural properties of certain crown decompositions potentially pave the way to achieving this goal.
Lastly, all missing details, i.e. omitted proofs and algorithms as well as more detailed explanations can be found in the appendix.

### v2

@@ Abstract
Over the past decades, various metrics have emerged in graph theory to grasp the complex nature of network vulnerability. In this paper, we study two specific measures: (weighted) vertex integrity (wVI) and (weighted) component order connectivity (wCOC). These measures not only evaluate the number of vertices that need to be removed to decompose a graph into fragments, but also take into account the size of the largest remaining component. The main focus of our paper is on kernelization algorithms tailored to both measures. We capitalize on the structural attributes inherent in different crown decompositions, strategically combining them to introduce novel kernelization algorithms that advance the current state of the field. In particular, we extend the scope of the balanced crown decomposition provided by Casel et al. [DBLP:conf/esa/Casel0INZ21] and expand the applicability of crown decomposition techniques.
In summary, we improve the vertex kernel of VI from $p^{3}$ to $3p^{2}$, and of wVI from $p^{3}$ to $3(p^{2}+p^{1.5}p_{\ell})$, where $p_{\ell}<p$ represents the weight of the heaviest component after removing a solution. For wCOC we improve the vertex kernel from $\mathcal{O}(k^{2}W+kW^{2})$ to $3\mu(k+\sqrt{\mu}W)$, where $\mu=\max(k,W)$. We also give a combinatorial algorithm that provides a $2kW$ vertex kernel in fixed-parameter tractable time when parameterized by $r$, where $r\leq k$ is the size of a maximum $(W+1)$-packing. We further show that the algorithm computing the $2kW$ vertex kernel for COC can be transformed into a polynomial algorithm for two special cases, namely when $W=1$, which corresponds to the well-known vertex cover problem, and for claw-free graphs. In particular, we show a new way to obtain a $2k$ vertex kernel (or to obtain a 2-approximation) for the vertex cover problem by only using crown structures.
@@ 1 Introduction
In the study of graph theory different scales have emerged over the past decades to capture the complex nature of network vulnerability. While the main focus is on the connectivity of vertices and edges, subtle aspects of vulnerability such as the number of resulting components, the size distribution of the remaining components, and the disparity between them are becoming increasingly interesting [DBLP:conf/esa/BentertHK23, DBLP:journals/tcs/GimaHKKO22, DBLP:conf/walcom/GimaHKM0O24, DBLP:journals/corr/abs-2402-09971, DBLP:journals/algorithmica/GimaO24, DBLP:conf/isaac/LampisM21]. Our focus in this study is on two specific ways to measure vulnerability: (weighted) vertex integrity and (weighted) component order connectivity. These measures not only evaluate the number of vertices that need to be removed to decompose a graph into fragments, but also take into account the size of the largest remaining component. Incorporating these aspects provides a more comprehensive understanding of network resilience.
Informally, vertex integrity (VI) is a model for the balance between removing few vertices and keeping small connected parts of a graph. More formally, given a graph $G=(V,E)$ and a number $p\in\mathds{N}$, the task for VI is to find a set of vertices $S\subseteq V$ such that $|S|$ plus the size of the largest component when removing $S$ from $G$ is at most $p$. In the vertex weighted version (wVI), the goal is bounding the total weight of the removed vertices plus the weight of the heaviest component by $p$. This problem was introduced by Barefoot et al. [barefoot1987vulnerability] as a way to measure vulnerability of communication networks. Recently, it has drawn the interest of the parameterized complexity community due to its status as a natural parameter that renders numerous NP-hard problems amenable to fixed parameter tractability (FPT). This means that for these problems, solutions can be computed within a time frame represented by $f(p)\cdot n^{\mathcal{O}(1)}$, where $f$ is a computable function [DBLP:conf/isaac/LampisM21]. It is interesting to see how vertex integrity relates to other well-known measures of network structure. It imposes greater constraints compared to metrics such as treedepth, treewidth, or pathwidth, as the vertex integrity of a graph serves as an upper bound for these parameters. However, it encompasses a wider range of scenarios compared to vertex cover, where a vertex cover of a graph is an upper bound for its vertex integrity. This makes it a key in understanding how to efficiently solve problems in the world of network analysis.
@@
The focus of the paper is on kernelization algorithms tailored for both weighted and unweighted versions of VI and COC when parameterized by $p$ and $k+W$, respectively. Kernelization algorithms can be thought of as formalized preprocessing techniques aimed to reduce insances of optimization problems. Of particular interest in this work are crown decompositions, which are generally used as established structures for safe instance reduction — where “safe” means that any optimal solution to the reduced instance can efficiently be transformed into an optimal solution of the original one. Essentially, a crown decomposition partitions the vertex set into three disjoint components: the crown, the head, and the body. Here, the head acts as a separator between the crown and the body. This structural arrangement becomes useful when specific relationships between the head and the crown are required. Such relationships ultimately enable us to shrink instances by eliminating these designated parts from the graph. The properties of this structural layout, coupled with its existence depending on the instance size, enable the development of efficient kernelization algorithms for different problem domains. Notably, crown decompositions have also recently found utility in approximation algorithms for graph packing and partitioning problems [DBLP:conf/esa/Casel0INZ21]. For further exploration of crown decompositions, including their variations and applications, we recommend the comprehensive survey paper by Jacob et al. [DBLP:journals/algorithms/JacobMR23].
@@
Recently, vertex integrity (VI) has received considerable attention as a structural graph parameter [DBLP:conf/esa/BentertHK23, DBLP:journals/tcs/GimaHKKO22, DBLP:journals/algorithmica/GimaO24, DBLP:conf/isaac/LampisM21]. Some results in the literature are stated in terms of the fracture number of the graph [DBLP:journals/algorithmica/GanianOR21], which is a closely related parameter that takes the maximum of the modulator size and the largest remaining component; in particular, fracture number is always within a factor of two from vertex integrity. Gima et al. [DBLP:conf/walcom/GimaHKM0O24] conducted a systematic investigation into structural parameterizations of computing the VI and wVI which was further extended by Hanaka et al. [DBLP:journals/corr/abs-2402-09971]. Additionally, there are notable results concerning special graph classes [DBLP:journals/dam/BaggaBGLP92, clark1987computational, DBLP:journals/algorithmica/DrangeDH16, DBLP:journals/dam/KratschKM97, DBLP:journals/ijcm/LiZZ08]. In our context, regarding related work on VI and wVI, Fellows and Stueckle presented an algorithm that solves the problem in time $\mathcal{O}(p^{3p}n)$ [fellows1989immersion]. This was subsequently improved by Drange et al. [DBLP:journals/algorithmica/DrangeDH16], even for the weighted case, to $\mathcal{O}(p^{p+1}n)$. In the same paper, they presented the first vertex-kernel of size $p^{3}$ for both VI and wVI.
@@
Considering, COC and wCOC, it is unlikely that kernelization algorithms for these problems can be achieved by considering $k$ or $W$ alone in polynomial time. Indeed, $W=1$ corresponds to the NP-hard vertex cover problem, which shows that $W$ (alone) is not a suitable parameter. For the parameter $k$, the problem is $W[1]$-hard even when restricted to split graphs [DBLP:journals/algorithmica/DrangeDH16]. These lower bounds lead to the study of parameterization by $k+W$. The best known algorithm with respect to these parameters finds a solution in time $n^{O(1)}\cdot 2^{\mathcal{O}(\log(W)\cdot k)}$ [DBLP:journals/algorithmica/DrangeDH16]. Unless the exponential time hypothesis fails, the authors prove that this running time is tight in the sense that there is no algorithm that solves the problem in time $n^{\mathcal{O}(1)}\cdot 2^{o(\log(W)\cdot k)}$. The best known approximation algorithm has a multiplicative gap guarantee of $\mathcal{O}(\log(W))$ to the optimal solution with a running time of $n^{\mathcal{O}(1)}\cdot 2^{\mathcal{O}(W)}$ [DBLP:journals/corr/Lee16c]. In [DBLP:journals/corr/Lee16c], the authors also showed that the superpolynomial dependence on $W$ may be needed to achieve a polylogarithmic approximation. Using this algorithm as a subroutine, the vertex integrity can be approximated within a factor of $\mathcal{O}(\log(\text{OPT}))$, where OPT is the vertex integrity.
@@
Regarding kernelization algorithms, there is a sequence of results which successively improve the vertex-kernel of COC. The first results came from Chen et al. [DBLP:journals/tcs/ChenFSWY19] and Drange et al. [DBLP:journals/algorithmica/DrangeDH16], who provided kernels of size in $\mathcal{O}(kW^{3})$ and $\mathcal{O}(k^{2}W+kW^{2})$, respectively. The result of Drange et al. also holds for wCOC and is the only result for this case. This was improved simultaneously by Xiao [DBLP:journals/jcss/Xiao17a] as well as by Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16] to a $\mathcal{O}(kW^{2})$ kernel. These works also provide the first $\mathcal{O}(kW)$ kernels, but with different constants and running times. Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16] present a $2kW$ kernel in a running time of $n^{\mathcal{O}(W)}$ by using linear programming (LP) methods with an exponential number of constraints. The runtime can be improved to $2^{\mathcal{O}(W)}\cdot n^{\mathcal{O}(1)}$ as already mentioned in the book of Fomin et al. [fomin2019kernelization] (Section 6.4.2). Roughly speaking, the idea is to use the ellipsoid method with separation oracles to solve the linear program, where the separation oracle uses a method called color coding to find violated constraints that makes it polynomial in $W$. Note that if $W$ is a constant this $2kW$ kernel is a polynomial time kernel improving on some previous results. This includes for instance the improvement of the $5k$ kernel provided by Xiao and Kou [DBLP:conf/tamc/XiaoK17] to a $4k$ kernel for the well-studied $P_{2}$-covering problem, where a $P_{2}$ is a path with 2 edges. The first linear kernel in both parameters in polynomial time, i.e. an $\mathcal{O}(kW)$ vertex kernel, is presented by Xiao [DBLP:journals/jcss/Xiao17a], who provides a $9kW$ vertex kernel. Finally, this was improved by Casel et al. [DBLP:conf/esa/Casel0INZ21] to a $3kW$ vertex kernel, which also holds for a more general setting. Namely, to find $k$ vertices in a vertex weighted graph such that after their removal each component weighs at most $W$. Note that the weights of the separator, i.e. the chosen $k$ vertices, play no role in this problem compared to wCOC. With the exception of the $\mathcal{O}(k^{2}W+kW^{2})$ vertex kernel of Drange et al. [DBLP:journals/algorithmica/DrangeDH16], all achieved vertex kernels essentially use crown structures.
Our Contribution
@@
The provided running times are based on an input graph $G=(V,E)$. We improve the vertex kernel for VI from $p^{3}$ to $3p^{2}$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$. For wVI, we improve it to $3(p^{2}+p^{1.5}p_{\ell})$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$, where $p_{\ell}\leq p$ represents the weight of the heaviest component after removing a solution.
@@
To better explain the results of COC and wCOC, consider the problem of a maximum $\lambda$-packing for $\lambda\in\mathds{N}$. Given a (vertex weighted) graph $G$, this problem aims to maximize the number of disjoint connected subgraphs, each of size (or weight) at least $\lambda$. It is worth noting that the size of a maximum $(W+1)$-packing serves as a lower bound on the size of an optimal solution of COC (or wCOC), since each element in the packing must contain at least one vertex of it — in terms of linear programming, it is the dual of COC.
@@
For wCOC we improve the vertex kernel of $\mathcal{O}(k^{2}W+kW^{2})$ to $3\mu(k+\sqrt{\mu}W)$ in time $\mathcal{O}\left(r^{2}k|V||E|\right)$, where $\mu=\max(k,W)$ and $r\leq|V|$ is the size of a maximum $(W+1)$-packing. For the unweighted version, we provide a $2kW$ vertex kernel in an FPT-runtime of $\mathcal{O}(r^{3}|V||E|\cdot r^{\min(3r,k)})$, where $r\leq k$ is the size of a maximum $(W+1)$-packing. Comparing this result with the state of the art, disregarding the FPT-runtime aspect, we improve upon the best-known polynomial algorithm, achieving a kernel of size $3kW$ [DBLP:conf/esa/Casel0INZ21]. A $2kW$ vertex kernel is also presented in [DBLP:conf/iwpec/KumarL16], albeit with an exponential runtime using linear programming methods, which, as mentioned, can be enhanced to an FPT-runtime regarding parameter $W$. In contrast, our result is entirely combinatorial and has an FPT-runtime in the parameter of a maximum $(W+1)$-packing $r\leq k$. It should be noted that, strictly speaking, the $2kW$ vertex kernel of [DBLP:conf/iwpec/KumarL16] and our work cannot be considered a kernel, given that the runtime dependency is exponential with respect to the parameters $W$ and $k$, respectively. However, for the sake of simplicity, we refer to it as a kernel, with an explicit mention of the runtime dependency. As previously stated, note that it is unlikely that an FPT-runtime will be sufficient for solving COC (or wCOC) when considering either $W$ or $k$ alone. We further show that the algorithm computing the $2kW$ vertex kernel for COC can be transformed into a polynomial algorithm for two special cases. The first case arises when $W=1$, i.e., for the vertex cover problem. Here, we provide a new method for obtaining a vertex kernel of $2k$ (or obtaining a $2$-approximation) using only crown decompositions. The second special case is for the restriction of COC to claw-free graphs. Notably, it remains open whether there is a polynomial-time algorithm for COC on claw-free graphs, or whether the problem is NP-hard on this class; we defer this question to future work.
@@
Regarding these special cases, until 2017, a $2k$ vertex kernel for $W=1$ was known through crown decompositions, albeit computed using both crown decompositions and linear programming. Previously, only a $3k$ vertex kernel was known using crown decompositions alone. In 2018, Li and Zhu [DBLP:journals/tcs/LiZ18] provided a $2k$ vertex kernel solely based on crown structures. They refined the classical crown decomposition, which possessed an additional property allowing the remaining vertices of the graph to be decomposed into a matching and odd cycles after exhaustively applying the corresponding reduction rule. In contrast, our algorithm identifies the reducible structures, which must exist if the size of the input graph exceeds $2k$. For claw-free graphs, we are not aware of previous work regarding kernels for VI or COC on this class. However, the classical result of Minty [DBLP:journals/jct/Minty80] states that one can find vertex cover of smallest weight in a claw-free graph in polynomial time.
Unfortunately, we were unable to transform the FPT-runtime kernelization algorithm into a polynomial-time algorithm in general. Nevertheless, we believe that our insights into the structural properties of crown decompositions pave the way to achieving this goal.